OpenStudy (meilendurcer):

Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

I think x would be -3 -3|x + 7| = -12 -3|-3 + 7| = -12 the absolute value of -3 +7 = 4 -3 x 4 = 12 ? correct me if i'm wrong lol

OpenStudy (meilendurcer):

@vt Yeah the only answers I have are" \[x= \frac{ 5 }{ 3 }, x= \frac{ 19 }{ 3 }\] x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (anonymous):

Oh....well then I have no clue haha sorry

OpenStudy (meilendurcer):

lol its ok thanks for helping though :D